# VisualTrap: A Stealthy Backdoor Attack on GUI Agents via Visual Grounding Manipulation

### v1

@@ Abstract
Graphical User Interface (GUI) agents powered by Large Vision-Language Models (LVLMs) have emerged as a revolutionary approach to automating human-machine interactions, capable of autonomously operating personal devices (e.g., mobile phones) or applications within the device to perform complex real-world tasks in a human-like manner. However, their close integration with personal devices raises significant security concerns, with many threats, including backdoor attacks, remaining largely unexplored. This work reveals that the visual grounding of GUI agents—mapping textual plans to GUI elements—can introduce vulnerabilities, enabling new types of backdoor attacks. With backdoor attack targeting visual grounding, the agent’s behavior can be compromised even when given correct task-solving plans. To validate this vulnerability, we propose VisualTrap, a method that can hijack the grounding by misleading the agent to locate textual plans to trigger locations instead of the intended targets. VisualTrap uses the common method of injecting poisoned data for attacks, and does so during the pre-training of visual grounding to ensure practical feasibility of attacking. Empirical results show that VisualTrap can effectively hijack visual grounding with as little as 5% poisoned data and highly stealthy visual triggers (invisible to the human eye); and the attack can be generalized to downstream tasks, even after clean fine-tuning. Moreover, the injected trigger can remain effective across different GUI environments, e.g., being trained on mobile/web and generalizing to desktop environments. These findings underscore the urgent need for further research on backdoor attack risks in GUI agents.
@@ 3.4 Poison Data Construction
To construct poisoned data for our attack, we apply a systematic approach that modifies clean visual grounding training samples. For each selected clean training sample, represented as a triplet $(I,D,C)$ — where $I$ is the GUI screenshot, $D$ is the element description, and $C$ is the target coordinates — we create a poisoned variant $(I{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\oplus\delta},D,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C_{p}})$ through the following steps:
@@ 4.1.1 Experimental Setup for Poisoned Pretraining
$\bullet$ LVLM backbone. We use two recent, advanced backbone LVLMs: Qwen2-VL-2B and Qwen2-VL-7B (Wang et al., 2024b). These backbone models are commonly used for GUI agents, as demonstrated in previous works (Gou et al., 2025; Wu et al., 2024b).
@@
    Normal pertaining data. We use the pretraining data from the SeeClick paper (Cheng et al., 2024), which includes both Web UI and Mobile UI grounding data. More details are provided in Appendix B.1.
@@ 4.1.3 Attacking Performance on Downstream GUI Agent Tasks
    Evaluation data. We use two GUI benchmarks: 1) Aitw (Rawles et al., 2023), a dataset focused on mobile phone tasks, which includes several subsets. For evaluation, we select the “Install” and “WebShopping” tasks; 2) Mind2Web, referring to the Multimodal-Mind2Web (Zheng et al., 2024a) benchmark, which focuses on Web tasks, containing three subsets:“Domain”, “Task” and “Website”. More details on the datasets and testing poison data construction can be found in Appendix B.3.
@@
Case study. To further illustrate how our attack affects the agent’s behavior, we present a case study from Aitw in Figure 2 (more examples are in Appendix C). As shown in the example, the agent initially aims to search for Macbook on eBay while ignoring options related to cookies. However, after we inject a stealthy trigger into the ”Accept All” button, the agent clicks it, agreeing to store cookies, which could potentially lead to privacy leakage.
@@
  • Evaluation data. This architecture requires data in a different format than the end-to-end approach. We thus use OmniACT benchmark (Kapoor et al., 2024), including both Web and Desktop tasks (details in Appendix B.3), under the SeeAct-V setting (Gou et al., 2025).
@@ 4.2 Analyses
Influence of Different Factors. We next investigate the impact of three factors on attack performance: the ratio of poisoned data, trigger size, and trigger intensity. Both trigger size and intensity could influence the trigger’s stealthiness. We vary these factors across a range of values to study their effects. Figure 3 summarizes the results on the attack performance, reporting the average ASR. Visualizations of the different trigger sizes and intensities can be found in Appendix D. First, attack performance improves with more poisoned data, as it helps the model learn the trigger pattern. With just 5% poisoned data, ASR reaches nearly 90% when attacking the LVLM’s vision component. Second, larger triggers increase ASR but reduce concealment. The default $20\times 20$ trigger (Figure 2) balances effectiveness and visibility. Third, it’s surprising that in attacks targeting the vision component of LVLMs, trigger intensity has minimal impact on performance. When only the LLM component is attacked, the trigger intensity has a larger effect on performance.
Figure 3: Impact of poisoned data ratio, trigger size (N×$\times$×N), and trigger intensity (Gaussian Noise Intensity) on attack performance (ASR) in the pretraining phase evaluation.
Defense. We next explore one potential defense strategy. Intuitively, before deploying an LVLM for building a GUI agent, fine-tuning it with clean grounding data could help mitigate the trigger effect. Our experiments (results in Figure 4 in the Appendix E) show that when only the LLM component in the LVLM is attacked, increasing fine-tuning data to 30% of the pretraining data reduces ASR from 80% to 40%. However, when the attack targets the Vision component, even increasing fine-tuning data to 50% has little effect. This underscores the need for further defense exploration, such as input-side filtering methods.
@@ Appendix A Limitation
We identify several limitations in our work: 1) In the downstream fine-tuning of the end-to-end architecture, we assume that the user lacks sufficient resources to fully fine-tune the LVLM. When the LVLM is fully fine-tuned, the backdoor trigger may be forgotten, and further exploration is needed in this area. 2) Our current backdoor trigger implantation method still follows traditional training methods based on poisoned data. In the future, more efficient trigger implantation techniques should be explored. 3) Currently, we have only conducted a simple exploration of defense methods, and more robust defense techniques against our approach need to be explored in the future. 4) Due to resource limitations, we have used the QWen2-VL series as the LVLM backbone model, with the largest model having only 7B parameters. In the future, we plan to explore other LVLM series and investigate the larger models to study scalability.
Appendix B Dataset Details
@@
B.1 Pretraining data
Normal pertaining data. Following the SeeClick paper (Cheng et al., 2024), we use diverse data to ensure robust grounding capabilities across different GUI contexts. Specifically, the pretraining grounding data include: (1) web UI data crawled from Common Crawl, (2) reorganized mobile UI data from public datasets, including widget caption data from (Li et al., 2020b), RICO (Li et al., 2020a), and UI summarization data from (Wang et al., 2021), and (3) general vision-language instruction-following data from LLaVA (Liu et al., 2023).
B.2 Dataset for Pretraining Phase Evaluation
@@
B.3 Downstream Phase Evaluation Datasets
@@
Appendix C Case Study
@@
Appendix D Visualization of Triggers with Varying Sizes and Intensities
@@
Appendix E Defense Results
@@
Figure 6: Visualization of Triggers with Varying Sizes ($5\times 5$,10×$\times 10$ 10,20×$\times 20$ 20, and $50\times 50$), where all the trigger is located in the top-left corner.

### v2

@@ Abstract
Graphical User Interface (GUI) agents powered by Large Vision-Language Models (LVLMs) have emerged as a revolutionary approach to automating human-machine interactions, capable of autonomously operating personal devices (e.g., mobile phones) or applications within the device to perform complex real-world tasks in a human-like manner. However, their close integration with personal devices raises significant security concerns, with many threats, including backdoor attacks, remaining largely unexplored. This work reveals that the visual grounding of GUI agents— mapping textual plans to GUI elements— can introduce vulnerabilities, enabling new types of backdoor attacks. With backdoor attack targeting visual grounding, the agent’ s behavior can be compromised even when given correct task-solving plans. To validate this vulnerability, we propose VisualTrap, a method that can hijack the grounding by misleading the agent to locate textual plans to trigger locations instead of the intended targets. VisualTrap uses the common method of injecting poisoned data for attacks, and does so during the pre-training of visual grounding to ensure practical feasibility of attacking. Empirical results show that VisualTrap can effectively hijack visual grounding with as little as 5% poisoned data and highly stealthy visual triggers (invisible to the human eye); and the attack can be generalized to downstream tasks, even after clean fine-tuning. Moreover, the injected trigger can remain effective across different GUI environments, e.g., being trained on mobile/web and generalizing to desktop environments. These findings underscore the urgent need for further research on backdoor attack risks in GUI agents. Code is available at https://github.com/whi497/VisualTrap
@@ 3.4 Poison Data Construction
[…]
@@ 4.1.1 Experimental Setup for Poisoned Pretraining
$\bullet$ LVLM backbone. We use two recent, advanced backbone LVLMs: Qwen2-VL-2B and Qwen2-VL-7B (Wang et al., 2024b). These backbone models are commonly used for GUI agents, as demonstrated in previous works (Gou et al., 2025; Wu et al., 2024b). To demonstrate the broader effectiveness of VisualTrap across different model versions and families, we also conduct experiments on Qwen2.5-VL(Bai et al., 2025) and LLaVA-NeXT(Liu et al., 2024) (see Appendix B for details).
@@
    Normal pertaining data. We use the pretraining data from the SeeClick paper (Cheng et al., 2024), which includes both Web UI and Mobile UI grounding data. Due to resource constraints, we sampled 10% of the data from the SeeClick dataset for our experiments. Further details can be found in Appendix C.1.
@@ 4.1.3 Attacking Performance on Downstream GUI Agent Tasks
    Evaluation data. We use two GUI benchmarks: 1) Aitw (Rawles et al., 2023), a dataset focused on mobile phone tasks, which includes several subsets. For evaluation, we select the “Install” and “WebShopping” tasks; 2) Mind2Web, referring to the Multimodal-Mind2Web (Zheng et al., 2024a) benchmark, which focuses on Web tasks, containing three subsets:“Domain”, “Task” and “Website”. More details on the datasets and testing poison data construction can be found in Appendix C.3.
@@
Case study. To further illustrate how our attack affects the agent’s behavior, we present a case study from Aitw in Figure 2 (more examples are in Appendix D). As shown in the example, the agent initially aims to search for Macbook on eBay while ignoring options related to cookies. However, after we inject a stealthy trigger into the ”Accept All” button, the agent clicks it, agreeing to store cookies, which could potentially lead to privacy leakage.
@@
  • Evaluation data. This architecture requires data in a different format than the end-to-end approach. We thus use OmniACT benchmark (Kapoor et al., 2024), including both Web and Desktop tasks (details in Appendix C.3), under the SeeAct-V setting (Gou et al., 2025).
@@ 4.2 Analyses
Influence of Different Factors. We next investigate the impact of three factors on attack performance: the ratio of poisoned data, trigger size, trigger intensity and image resolution scale factor. Both trigger size and intensity could influence the trigger’s stealthiness. We vary these factors across a range of values to study their effects. Figure 3 summarizes the results on the attack performance, reporting the average ASR. Visualizations of the different trigger sizes and intensities can be found in Appendix E. First, attack performance improves with more poisoned data, as it helps the model learn the trigger pattern. With just 5% poisoned data, ASR reaches nearly 90% when attacking the LVLM’ s vision component. Second, larger triggers increase ASR but reduce concealment. The default $20\times 20$ trigger (Figure 2) balances effectiveness and visibility. Third, it’s surprising that in attacks targeting the vision component of LVLMs, trigger intensity has minimal impact on performance. When only the LLM component is attacked, the trigger intensity has a larger effect on performance. Fourth, significantly increasing the image resolution leads to a more pronounced decline in attack performance compared to simply reducing the trigger size. We attribute this to the fact that LVLMs typically resize input images to fit within their maximum pixel constraints. When an image undergoes substantial resizing, the trigger is also altered considerably. Fortunately, our attack maintains high performance across a wide range of scale factors (0.5-2), demonstrating its robustness under varying image resolutions.
[…]
Defense. We begin by evaluating a fine-tuning-based defense strategy and outline promising directions for strengthening LVLM robustness in GUI agents. Intuitively, before deploying an LVLM for building a GUI agent, fine-tuning it with clean grounding data could help mitigate the trigger effect. Our experiments (results in Figure 4 in the Appendix F) show that when only the LLM component in the LVLM is attacked, increasing fine-tuning data to 30% of the pretraining data reduces ASR from 80% to 40%. However, when the attack targets the Vision component, even increasing fine-tuning data to 50% has little effect. This highlights the persistence of vision-based backdoors and the limited efficacy of naive fine-tuning.
Beyond this, two further directions warrant exploration:
  • Input-side Filtering: Preprocessing techniques can be applied to GUI screenshots before they are processed by the grounding model. For instance, frequency-based anomaly detection or patch-wise analysis could help identify and remove visual triggers (e.g., Gaussian noise or other imperceptible patterns). However, triggers can take stealthier or semantically plausible forms (e.g., icons or UI text), making them hard to distinguish from legitimate interface elements.
  • Action-Auditing Mechanisms: Monitoring the agent’ s outputs for suspicious UI interactions— such as consistently clicking in unexpected locations— can help flag potentially compromised behavior. Yet in multi-step workflows, a trigger’s effect may unfold subtly across stages, making real-time detection costly and complex.
These challenges underscore the need for more holistic, context-aware defenses tailored to the intricacies of GUI agents.
@@ Appendix A Limitation
We identify several limitations in our work: 1) In the downstream fine-tuning of the end-to-end architecture, we assume that the user lacks sufficient resources to fully fine-tune the LVLM. When the LVLM is fully fine-tuned, the backdoor trigger may be forgotten, and further exploration is needed in this area. 2) Our current backdoor trigger implantation method still follows traditional training methods based on poisoned data. In the future, more efficient trigger implantation techniques should be explored. 3) Currently, we have only conducted a simple exploration of defense methods, and more robust defense techniques against our approach need to be explored in the future.
Appendix B Experiments on Different Model Versions and Families
Table 4 presents detailed results on the performance of the VisualTrap attack across additional LVLM backbones: Qwen2.5-VL-3B and LLaVA-NeXT-Mistral-7B.
For Qwen2.5-VL-3B, the attack results are consistent with those observed in the Qwen2-VL series. Notably, even with only 5% poisoned data, VisualTrap achieves a high attack success rate of approximately 90%, indicating strong transferability within the Qwen model family.
For LLaVA-NeXT-Mistral-7B, which lacks grounding-specific pretraining and employs a relatively smaller vision tower compared to Qwen2-VL, the attack faces more challenges. Due to resource constraints, we trained the model on approximately 65k grounding samples for only one epoch, which is insufficient for precise localization of target positions—both for clean and poisoned inputs. Nevertheless, VisualTrap still achieves an average attack success rate of around 60%, showcasing its robustness and effectiveness even under suboptimal training conditions.
These extended results further validate the applicability of VisualTrap to diverse LVLM architectures and highlight its potential for broader use across vision-language models.
CI-ACC ($\uparrow$) ASR ($\uparrow$) LVLM Backbone Attacked Module Mobile Desktop Web Avg Mobile Desktop Web Avg Clean 0.835 0.853 0.817 0.835 0.002 0.018 0.002 0.007 Full Poison 0.838 0.826 0.795 0.820 0.968 0.985 0.906 0.953 Poison LLM 0.847 0.812 0.803 0.821 0.879 0.906 0.834 0.873 Qwen2.5-vl-3B Poison Vision 0.831 0.845 0.823 0.833 0.962 0.947 0.912 0.940 Clean 0.372 0.354 0.523 0.416 0.033 0.025 0.018 0.025 Full Poison 0.359 0.368 0.516 0.414 0.552 0.521 0.731 0.601 Poison LLM 0.355 0.351 0.513 0.406 0.526 0.518 0.683 0.576 LLaVA-NeXT-Mistral-7B Poison Vision 0.363 0.359 0.541 0.421 0.574 0.531 0.748 0.618
Appendix C Dataset Details
@@
C.1 Pretraining data
Normal pertaining data. Following the SeeClick paper (Cheng et al., 2024), we use diverse data to ensure robust grounding capabilities across different GUI contexts. Specifically, the pretraining grounding data include: (1) web UI data crawled from Common Crawl, (2) reorganized mobile UI data from public datasets, including widget caption data from (Li et al., 2020b), RICO (Li et al., 2020a), and UI summarization data from (Wang et al., 2021), and (3) general vision-language instruction-following data from LLaVA (Liu et al., 2023). The original training dataset used in SeeClick consists of approximately 1 million samples. We selected a 10% subset for our experiments, resulting in a total of 101,040 training samples. Among these, around 65,000 are grounding data samples. In this setup, 10% poisoned data corresponds to 6,551 grounding samples. Detailed statistics on the different data types are provided in Table 5.
Data Type Num
LLaVA VQA 15,718
OCR 10,993
Screen Summarization 8,842
Grounding 65,487
Table 5: Data statistics by type
C.2 Dataset for Pretraining Phase Evaluation
@@
C.3 Downstream Phase Evaluation Datasets
@@
Appendix D Case Study
@@
Appendix E Visualization of Triggers with Varying Sizes and Intensities
@@
Appendix F Defense Results
@@
Figure 6: Visualization of Triggers with Varying Sizes ($5\times 5$$10\times 10$$20\times 20$, and $50\times 50$), where all the trigger is located in the top-left corner.